# Are Large Language Models Rational Investors?

### v1

@@ F.3 Risk-Preference Questionnaire Template
Appendix G Result
G.1 Analysis of Direct News Events
The examination of direct news events involves a detailed statistical analysis of the responses generated by various Large Language Models (LLMs) to specific news items. This analysis primarily focuses on the distribution of scores assigned by LLMs to each news event, encompassing key statistical measures such as the mean, variance, highest, and lowest scores. Such an approach is instrumental in assessing the consistency and rationality of LLMs’ interpretations of financial news.
@@ G.1 Analysis of Direct News Events
In parallel with the examination of model biases, our study also delves into the temporal evolution of Large Language Models within distinct series, attributing changes to factors such as model size or software updates. To this end, we employ box line comparison charts as a visual tool to elucidate the developmental trajectories of models within the Baichuan, GLM, and LLaMA series. These charts serve to highlight variations in model performance or bias over time, providing a clear visual representation of progression or shifts in model behavior. The comparative analyses for the Baichuan series, GLM series, and LLaMA series are depicted in Figure 30, Figure 31, and Figure 32, respectively. Through these visual comparisons, we aim to offer insights into how advancements or modifications in model architecture and capabilities influence their analytical outcomes and biases.
G.2 Analysis of COT News
@@ G.3 Analysis of Interactions
G.4 Analysis of Risk-preference Questions
@@
Subsequently, our focus shifted to the models’ adherence to explicit instructions regarding risk aversion. By inputting the directive "You are a risk averse person," we were able to quantify each model’s compliance through the risk averse ratio, the details of which are encapsulated in Table 7. This aspect of the study provides insight into the models’ capacity for context-based adaptability and their interpretation of subjective instructions.
Further, to ascertain the impact of the framing effect on model responses, a set of questions was translated to examine any discrepancies arising from linguistic variations. The findings from this segment of the study, highlighting the influence of translation on model outputs, are presented in Table 8. This analysis contributes to understanding the potential for framework effects to skew model perception and decision-making processes.
Lastly, we delved into the models’ susceptibility to loss aversion by introducing scenarios framed around loss. The extent of loss aversion bias manifesting in the models’ responses was rigorously analyzed, with the summarized results being showcased in Table 9. Through this comprehensive approach, we aim to unveil the multifaceted nature of biases in LLMs, particularly in the context of risk assessment and decision-making under uncertainty.
@@
| Model | Input Form | Risk Averter | Risk Neutral | Risk Lover |

### v2

@@ F.3 Risk-Preference Questionnaire Template
Appendix G FinCausal Dataset
In this section, we introduce the construction process of the FinCausal dataset, which requires the acquisition of relevant causal knowledge from past financial text materials. This process can be mainly divided into data collection, deduplication, segmentation, and knowledge extraction. The main flowchart is shown in Figure 11.
Firstly, we crawled 500,000 research reports from the internet, ranging from 2020 to 2023, including individual stock research, industry analysis, and macro analysis. We used regular matching and the FastText language filter for classification, mainly retaining Chinese A-Share individual stock research and industry analysis. Since individual stock reports rarely describe causal relationships for negative events, we further crawled some news analyses and stock forum comments to enrich the description of individual stock causal knowledge.
After filtering the research report data, we used the MinHash algorithm to perform document-level deduplication on all content. To extract sentences with causal expressions, we meticulously categorized the content of the research reports into seven distinct types, which included ordinary sentences, causal sentences, news-related content, recommendation ratings, investment advice, risk warnings, and researcher information. We then manually annotated a comprehensive dataset comprising 3,000 pieces of data to train a sophisticated BGE+TextCNN classification model. This model was specifically designed to discern and categorize the various types of sentences present in the financial reports, with a particular focus on identifying those that convey causal relationships.
For each piece of extracted knowledge from a research report or comment, we concatenated one sentence before and after it into a paragraph. Subsequently, we aggregated all relevant paragraphs from a report to form a comprehensive context. Utilizing our meticulously chosen LLM, we conducted causal knowledge extraction from these contexts. Through this process, we successfully obtained 200,000 pieces of industry causal knowledge and 2,000 pieces of individual stock causal knowledge, thereby enriching our dataset with valuable insights into the causal relationships within the financial domain. Here are some examples of FinCausal:
The company may consider conducting targeted issuance in order to expand its business scale, make capital expenditures, or invest in research and development.
During the epidemic, the demand for remote work and online collaboration has increased, driving the development of related software service companies.
Appendix H Result
H.1 Analysis of Direct News Events
The examination of news events involves a detailed statistical analysis of the responses generated by various Large Language Models (LLMs) to specific news items. This analysis primarily focuses on the distribution of scores assigned by LLMs to each news event, encompassing key statistical measures such as the mean, variance, highest, and lowest scores. Such an approach is instrumental in assessing the consistency and rationality of LLMs’ interpretations of financial news.
@@ H.1 Analysis of Direct News Events
In parallel with the examination of model biases, our study also delves into the temporal evolution of Large Language Models within distinct family, attributing changes to factors such as model size or software updates. To this end, we employ box line comparison charts as a visual tool to elucidate the developmental trajectories of models within the Baichuan, GLM, and LLaMA family. These charts serve to highlight variations in model performance or bias over time, providing a clear visual representation of progression or shifts in model behavior. The comparative analyses for the Baichuan family, GLM family, and LLaMA family are depicted in Figure 33, Figure 34, and Figure 35, respectively. Through these visual comparisons, we aim to offer insights into how advancements or modifications in model architecture and capabilities influence their analytical outcomes and biases.
H.2 Analysis of COT News
@@ H.3 Analysis of Interactions
H.4 Analysis of Risk-preference Questions
@@
Subsequently, our focus shifted to the models’ adherence to explicit instructions regarding risk aversion. By inputting the directive "You are a risk averse person," we were able to quantify each model’s compliance through the risk averse ratio, the details of which are encapsulated in LABEL:tab:instruct_risk. This aspect of the study provides insight into the models’ capacity for context-based adaptability and their interpretation of subjective instructions.
Further, to ascertain the impact of the framing effect on model responses, a set of questions was translated to examine any discrepancies arising from linguistic variations. The findings from this segment of the study, highlighting the influence of translation on model outputs, are presented in LABEL:tab:translation_difference. This analysis contributes to understanding the potential for framework effects to skew model perception and decision-making processes.
Lastly, we delved into the models’ susceptibility to loss aversion by introducing scenarios framed around loss. The extent of loss aversion bias manifesting in the models’ responses was rigorously analyzed, with the summarized results being showcased in LABEL:tab:Loss_Aversion_Bias. Through this comprehensive approach, we aim to unveil the multifaceted nature of biases in LLMs, particularly in the context of risk assessment and decision-making under uncertainty.
@@
| Model | Method | Risk Averter | Risk Neutral | Risk Lover |